# Critical dynamics and its interferometry in the one-dimensional p-wave-paired Aubry-André-Harper model

### v1

@@ I Introduction
The renowned Kibble-Zurek mechanism (KZM) was first proposed to understand the signatures of phase transitions that have occurred in the early universe, determining the density of defects left in the broken symmetry phase [1, 2]. Soon later, similar critical dynamics in condensed matter systems is uncovered to be described by KZM also [3]. Specifically, for a linear quench protocol, the distance from the critical point can be parameterized by $\epsilon(t)=(t-t_{c})/\tau_{Q}$, where $\tau_{Q}$ is the quench time and $t_{c}$ marks the time when the critical point is crossed. After crossing the critical point, the final density of defects exhibits the KZ scaling law, $\sim\tau_{Q}^{-dv/1+vz}$, where $d$ is the dimension of space, $\nu$ and $z$ are critical exponents. Since then, progresses have been achieved in both theoretical and experimental researches [4, 5, 6, 7, 8, 9]. A recent study revealed an interference effect of two successive critical dynamics realized by a properly designed quench protocol, which can be easily observed through the oscillatory KZ behavior in the defect density [10]. For fast quench rates, the KZ scaling (KZS) law will be broken. However, extensive studies have revealed that there are saturated and presaturated regimes beyond the KZ regime [5, 11, 12, 13, 14, 15, 16, 17, 18, 19].
On the other hand, the quantum localization was initially found in a randomly disordered system [20]. Since then, the research on the topic of localized-delocalized phase transition has attracted a lot of attention [21, 22, 23, 24, 25, 26, 27]. Later, a similar localized-delocalized phase transition was found to be realized in a quasiperiodic disordered system, the so-called Aubry-André-Harper (AAH) model for one-dimensional wires [28, 29]. It has stimulated a lot of theoretical and experimental researches on quasiperiodic systems, in which intriguing phenomena such as fractal spectrum, many-body localization, and mobility edge are revealed [30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49]. The critical dynamics of the AAH model have also been extensively studied, which shows a significant difference from that of the randomly disordered systems [50].
@@
The rest of the paper is organized as follows: In II, we briefly introduce the model Hamiltonian. In III.1, we demonstrate the two-plateau structure of the dynamical exponent. In IV, we employ two different quench dynamics, the one-way and round-trip quench protocol, to unravel the impact of the dynamical exponent on the critical dynamics . In IV.3, we study the quench dynamics by varying p-wave superconductivity pairing amplitude. At last, we give a summary.
Figure 1: The energy gap behavior and the quench protocols at the p-wave paring amplitude $\Delta=1$. The energy gap $\Delta_{g}$ as a function of potential strength $V$ by varying the size of system $L$. In the localized phase, the gap gradually closes with the increasing $L$. Moreover, two quench protocols are marked by different arrows. The one-way quench protocol linearly drives the system from the localized phase to the critical phase. The round-trip one consists of two successive linear ramps with the turning point $V_{rt}$, $\tau_{Q}$ and $\tau_{Q}^{\prime}$ denote quench times of two linear ramps respectively.
@@ II Model Hamiltonian
The phase diagram of the AAH model with p-wave superconducting pairing amplitude $\Delta$ and the quasi-periodic potential strength $V$ has been thoroughly investigated in previous research [51, 52, 55]. Three phases, named extended, critical, and localized, are explored. The system undergoes the extended-critical phase transition at $V=\left|V-\Delta\right|$ and the critical-localized phase transition at $V=\left|V+\Delta\right|$.
Figure 2: The dynamical exponent in the gapless localized phase. (a) The dynamical exponent $z$ as a function of the distance from the critical point $\epsilon$. Two plateaus at $z\approx 1$, and $1.388$, emerge at far and near from the critical point ($\epsilon=0$). The fitted straight lines for each subsequences gives: , (c) $\Delta_{g}\sim L^{1.3872\pm 0.0003}$ and $L^{1.3877\pm 0.0005}$ with $\epsilon=10^{-4}$, (b) $\Delta_{g}\sim L^{1.28\pm 0.04}$ and $L^{1.17\pm 0.06}$ with $\epsilon=10^{-2}$, and (d) we obtained $\Delta_{g}\sim L^{0.999\pm 0.0005}$ with $\epsilon=1$. These conclusions were derived from analyzing 500 different values of $\phi$.
@@ III.1 Dynamical exponent at $\Delta=1$
We set the p-wave pairing amplitude as $\Delta=1$ and the hopping amplitude $J=1$ to investigate the dynamical exponent for the critical-localized phase transition. In this case, the critical point of this phase transition is located at $V_{c}=J+\Delta=2$. Interestingly, the localized phase is found to be gapless within a large potential strength $V>V_{c}$(Fig. 1), which is characterized as a gap
@@
Figure 2(a) shows how the dynamical exponent $z$ depends on the distance from the critical point $\epsilon$ in the gapless localized phase. The exponent $z$ is determined by averaging the linear fits of the scaling relation $\Delta_{g}\sim L^{-z}$ across three Fibonacci subsequences: $F_{3l}$, $F_{3l+1}$, and $F_{3l+2}$, following the methodology established in Refs. [61, 58]. In Fig. 2(a), the plateau $z_{\text{I}}\approx 1$ holds with large $\epsilon$ and then gradually rises, eventually saturating at $z_{\text{II}}\approx 1.388$ as $\epsilon$ narrow, aligning with the dynamical exponent obtained at the critical point. [55, 57]. The dependence of the dynamical exponent on the potential strength in the gapless localized phase is similar to that of the random disorder system [62, 60]. However, unlike the random disorder system, $z$ does not diverge at the critical point but instead saturates to a constant. The gap as a function of the size of the system $L$ by varying the distance from the critical point is shown in Fig. 2(b), (c), and (d). In Fig. 2(b), near the critical point, dynamical exponents $z=1.3872\pm 0.0003$ and $1.3875\pm 0.0005$ are obtained from different subsequences with a negligible difference. However, this difference becomes inescapable, $z=1.28\pm 0.04$ and $1.17\pm 0.06$, with the distance of the critical point gradually increasing, see Fig. 2(c). When the potential strength is further away from the phase transition point, different subsequences will share a common power-law exponent $z=1$ (Fig. 2(d)).
@@ III.2 Dynamical exponent for varying $\Delta$
The dynamical exponent as a function of the reduced potential strength is shown in Fig. 3(b), where the curves with different p-wave superconductivity pairing amplitude $\Delta$ collapse into a universal scaling. Two plateaus are observed in the $z(\delta)$ profile, one near and the other far from the critical point.
@@ IV.1 One-way quench protocol
The one-way quench protocol is chosen to be inquired first. In this protocol, the reduced potential strength can be parameterized as follows:
@@
where the dynamical exponent $z(\hat{\epsilon})$ is obtained from (6). At $\epsilon=-\hat{\epsilon}$, the ground state gets excited and freeze-out with a corresponding correlation length $\hat{\xi}\sim\hat{\epsilon}^{-\nu}$. The state does not change until $t=\hat{t}$ when the relaxation time keeps pace with the evolving parameter successfully. After $\hat{t}$, the system gets adiabatic evolution, and the correlation length corresponding to $-\hat{t}$ is expected to populate. Consequently, the density of excitation probability for a one-dimensional system scales as follows:
@@
where $F$ is a non-universal scaling function[64, 65, 66]. Consequently, the scaled excitation density has a scaling function dependent on the scaled time $t/\hat{t}$. In conclusion, the information of the changing dynamical exponent is expected to be found in this KZ scaling law.
[…]
$\begin{split}n_{\text{ex}}&=n_{su}-CV_{i}^{3}\tau_{Q}^{2},\\ n_{\text{su}}&=\frac{1}{2}-\frac{\alpha}{V_{i}},\end{split}$ (15)
where C is a constant. When the quench rate declines and reaches the tuning point, $\tau_{Q}=\tau_{Q}^{S}$, the pre-saturate regime(PSR) appears because the initial potential strength is out of the impulse regime [19]. The defect’s density within the PSR is a linear combination of $\tau_{Q}^{1/2}$ and $\tau_{Q}^{3/2}$:
@@
Figure 4: The defects as a function of quench time $\tau_{Q}$ with system size $L=987$. The final defects are obtained from the linear quench by adjusting the quench rate from slow to very fast. After the fast quench regimes, the fit gives $n_{\text{ex}}\sim\tau_{Q}^{-0.507\pm 0.007}$ (KZ-I) and $n_{\text{ex}}\sim\tau_{Q}^{-0.420\pm 0.008}$ (KZ-II), corresponding to the fast and slow quench rate. The alternation of the KZ exponents reflects the two plateaus of the dynamical exponent $z\approx 1$ and $z\approx 1.388$. At saturate regime (SR), the density of defects is saturated to a plateau at first and then decreases with a scaling $n_{\text{ex}}\sim-\tau_{Q}^{2}$. In the pre-saturate regime (PSR), $n_{\text{ex}}$ declines as $1/2-0.52\tau_{Q}^{0.49}$ until $\tau_{Q}>1$.
@@
IV.1.1 Defect behavior after one-way quenching
@@
The final defects density $n_{\text{ex}}$ as a function of the quench time $\tau_{Q}$ is shown in Fig. 4 from slow to very fast quench rate with the initial potential strength $V_{i}=4$. According to the behavior of the defect density, we indicate the saturated, pre-saturated, and Kibble-Zurek regimes. In the SR, the saturate plateau $n_{\text{su}}=0.3865$ appears when the quench rate is very fast, as the dashed red line. A fitting of $(n_{\text{su}}-n_{\text{ex}})\sim\tau_{Q}^{s}$ yields $s=2.004\pm 0.005$, which is in good agreement with Eq. (IV.1.1). With the increasing of quench time, the initial potential strength $V_{i}$ is out of the non-adiabatic regime, $V_{i}>V+\hat{\epsilon}$, and quench dynamics turns into PSR. In this case, $n_{\text{ex}}$ diminishes as the expression $1/2-0.52\tau_{Q}^{0.49}$, which is revealed by a fitting of $(1/2-n_{\text{ex}})=\tau_{Q}^{s}$ yields $s=0.490\pm 0.009$. The $\tau_{Q}^{3/2}$ term in Eq.(16) does not appear because of the small fitting parameter $B$. After the quench time gets to $\tau_{Q}\approx 1$, the adiabatic-impulse-adiabatic hypothesis kicks in, and the KZ scaling law can reasonably predict the behavior of defects’ density.
@@
In the KZ regime, the power-law fit and Eq. (13) imply two dynamical exponents $z_{\text{I}}\approx 0.972\pm 0.027$ and $z_{\text{II}}\approx 1.381\pm 0.045$ for exact $\nu=1$, corresponding to the moderately fast and slow quench rates. The two values of dynamical exponents match the ones obtained by the finite-size scaling as shown in Fig. (2), $z=1$ and $1.388$, with systematic errors $3\%$ and $0.6\%$ respectively. In the moderately fast quench rate, $z_{\text{I}}$ closely resembles that of the transverse field Ising model. In the slow quench regime, the KZ exponent shifts to $z_{\text{II}}$, which is in good agreement with the known result [55]. ̵Specifically, for the moderately fast quench rate, the impulse regime is far from the critical point but does not extend beyond $V_{i}$ with the dynamical exponent staying at a constant $z(\hat{\epsilon}\gtrsim 1)\approx 1$. In this case, the defect’s density behaves as a power-law scaling $n_{\text{ex}}\sim\tau_{Q}^{-0.5}$. With the quench rate slowing down, the non-adiabatic regime becomes narrow and near the critical point. The dynamical exponent will saturate to a plateau $z(\hat{\epsilon}\rightarrow 0)\approx 1.388$, corresponding to the excitation probability scaling as $n_{\text{ex}}\sim\tau_{Q}^{-0.42}$, which in good agreement with the result in the past research[57]. This variable KZ exponent is similar to that in the random Ising chain affected by the Griffith phase [62, 60]. However, the KZ exponent will be steady because the dynamical exponent saturates to a constant at the critical point instead of diverging.
@@
The influence of the dynamical exponent is also reflected in the non-universal function (14), which discloses a universal behavior of the excitation probability in the impulse region. The defect’s density as a function of scaled time in the non-adiabatic region is shown in Fig. 5 (a) and (c) by varying quench time $\tau_{Q}$ belonging to KZ-I and KZ-II, respectively. Different scaling factors, $\tau_{Q}^{0.5}$ and $\tau_{Q}^{0.42}$ corresponding to the inverse of $\hat{n}_{\text{ex}}$, are multiplied in $n_{\text{ex}}$ to disclosed the non-universal function (14) in Fig. 5 (b) and (c).

### v2

@@ I Introduction
The renowned Kibble-Zurek mechanism (KZM) was first proposed to understand the signatures of phase transitions that have occurred in the early universe, determining the density of defects left in the broken symmetry phase [1, 2]. Soon later, similar critical dynamics in condensed matter systems is uncovered to be described by KZM also [3]. Specifically, for a linear quench protocol, the distance from the critical point can be parameterized by $\epsilon(t)=(t-t_{c})/\tau_{Q}$, where $\tau_{Q}$ is the quench time and $t_{c}$ marks the time when the critical point is crossed. After crossing the critical point, the final density of defects exhibits the KZ scaling law, $\sim\tau_{Q}^{-dv/1+vz}$, where $d$ is the dimension of space, $\nu$ and $z$ are critical exponents. Since then, progresses have been achieved in both theoretical and experimental researches [4, 5, 6, 7, 8, 9, 10, 11, 12, 13]. A recent study revealed an interference effect of two successive critical dynamics realized by a properly designed quench protocol, which can be easily observed through the oscillatory KZ behavior in the defect density [14]. For fast quench rates, the KZ scaling law will be broken. However, extensive studies have revealed that there are saturated and presaturated regimes beyond the KZ regime [15, 16, 17, 18, 19, 20, 21, 22, 23, 24].
On the other hand, the quantum localization was initially found in a randomly disordered system [25]. Since then, the research on the topic of localized-delocalized phase transition has attracted a lot of attention [26, 27, 28, 29, 30, 31, 32, 33]. Later, a similar localized-delocalized phase transition was found to be realized in a quasiperiodic disordered system, the so-called Aubry-André-Harper (AAH) model for one-dimensional wires [34, 35]. It has stimulated a lot of theoretical and experimental researches on quasiperiodic systems, in which intriguing phenomena such as fractal spectrum, many-body localization, and mobility edge are revealed [36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57]. The critical dynamics of the AAH model have also been extensively studied, which shows a significant difference from that of the randomly disordered systems [58, 59, 60].
@@
The rest of the paper is organized as follows: In II, we briefly introduce the model Hamiltonian. In III, we demonstrate the two-plateau structure of the dynamical exponent. In IV, we employ two different quench dynamics, the one-way and round-trip quench protocol, to unravel the impact of the dynamical exponent on the critical dynamics. At last, we give a summary.
Figure 1: The energy gap behavior and the quench protocols at the p-wave paring amplitude $\Delta=1$. The energy gap $\Delta_{g}$ is collected as a function of potential strength $V$ by varying the size of system $L$. In the localized phase, the gap gradually closes with the increasing $L$. Moreover, two quench protocols are marked by different arrows. The one-way quench protocol linearly drives the system from the localized phase to the critical phase. The round-trip one consists of two successive linear ramps with the turning point $V_{rt}$, $\tau_{Q}$ and $\tau_{Q}^{\prime}$ denote quench times of two linear ramps respectively.
@@ II Model Hamiltonian
The phase diagram of the AAH model with p-wave superconducting pairing amplitude $\Delta$ and the quasi-periodic potential strength $V$ has been thoroughly investigated in previous research [61, 62, 65]. Three phases, named extended, critical, and localized, are explored. The system undergoes the extended-critical phase transition at $V=\left|J-\Delta\right|$ and the critical-localized phase transition at $V=\left|J+\Delta\right|$.
Figure 2: The dynamical exponent in the gapless localized phase. (a) The dynamical exponent $z$ as a function of the distance from the critical point $\epsilon$. Two plateaus at $z\approx 1$, and $1.388$, emerge at far and near from the critical point ($\epsilon=0$). The fitted straight lines for each subsequences gives: , (b) $\Delta_{g}\sim L^{1.3872\pm 0.0003}$ and $L^{1.3877\pm 0.0005}$ with $\epsilon=10^{-4}$, (c) $\Delta_{g}\sim L^{1.28\pm 0.04}$ and $L^{1.17\pm 0.06}$ with $\epsilon=10^{-2}$, and (d) we obtained $\Delta_{g}\sim L^{0.999\pm 0.0005}$ with $\epsilon=1$. These conclusions were derived from analyzing 500 different values of $\phi$.
@@ III.1 Dynamical exponent at $\Delta=1$
We set the p-wave pairing amplitude $\Delta=1$ and the hopping amplitude $J=1$ to investigate the dynamical exponent for the critical-localized phase transition. In this case, the critical point of this phase transition is located at $V_{c}=\left|J+\Delta\right|=2$. Interestingly, the localized phase is found to be gapless within a large potential strength $V>V_{c}$ (Fig. 1), which is characterized as a gap
@@
Figure 2(a) shows how the dynamical exponent $z$ depends on the distance from the critical point $\epsilon$ in the gapless localized phase. The exponent $z$ is determined by averaging the linear fits of the scaling relation $\Delta_{g}\sim L^{-z}$ across three Fibonacci subsequences: $F_{3l}$, $F_{3l+1}$, and $F_{3l+2}$, following the methodology established in Refs. [71, 68]. In Fig. 2(a), the plateau $z_{\text{I}}\approx 1$ holds with large $\epsilon$ and then gradually rises, eventually saturating at $z_{\text{II}}\approx 1.388$ as $\epsilon$ narrow, aligning with the dynamical exponent obtained at the critical point [65, 67]. The dependence of the dynamical exponent on the potential strength in the gapless localized phase is similar to that of the random disorder system [72, 70]. However, unlike the random disorder system, $z$ does not diverge at the critical point but instead saturates to a constant. The gap as a function of the size of the system $L$ by varying the distance from the critical point is shown in Fig. 2(b), (c), and (d). In Fig. 2(b), near the critical point, dynamical exponents $z=1.3872\pm 0.0003$ and $1.3875\pm 0.0005$ are obtained from different subsequences with a negligible difference. However, this difference becomes inescapable, $z=1.28\pm 0.04$ and $1.17\pm 0.06$, with the distance of the critical point gradually increasing, see Fig. 2(c). When the potential strength is further away from the phase transition point, different subsequences will share a common power-law exponent $z=1$ (Fig. 2(d)).
@@ III.2 Dynamical exponent for varying $\Delta$
The dynamical exponent as a function of the reduced potential strength is shown in Fig. 3(b), where the curves with different $\Delta$ collapse into a universal scaling. Two plateaus are observed in the $z(\delta)$ profile, one near and the other far from the critical point.
@@ IV.1 One-way quench protocol
The one-way quench protocol is chosen to be inquired first. In this protocol, the distance from the critical point can be parameterized as follows:
@@
where the dynamical exponent $z(\hat{\epsilon})$ is obtained from Eq. (6). At $\epsilon=-\hat{\epsilon}$, the ground state gets excited and freeze-out with a corresponding correlation length $\hat{\xi}\sim\hat{\epsilon}^{-\nu}$. The state does not change until $t=\hat{t}$ when the relaxation time keeps pace with the evolving parameter successfully. After $\hat{t}$, the system gets adiabatic evolution, and the correlation length corresponding to $-\hat{t}$ is expected to populate. Consequently, the density of excitation probability for a one-dimensional system scales as follows:
@@
where $F$ is a non-universal scaling function [75, 76, 77]. Consequently, the scaled excitation density has a scaling function dependent on the scaled time $t/\hat{t}$.
However, as the quench speed gradually increases, the KZ scaling law gradually fails because of the adiabatic-impulse-adiabatic approximation failure. In the fast quench, the defect behavior can be divided into saturated and pre-saturated regimes. The saturated regime (SR) is defined by an extremely fast quench rate, where the KZ scaling law breaks down. In this regime, the defect density forms a plateau with $n_{\text{ex}}=n_{\text{su}}$. This plateau is attributed to sudden quench and depends on the initial potential strength $V_{i}$ [20]. For the transverse Ising model, the defect density within the SR scales as a function of $\tau_{Q}^{2}$ :
$\begin{split}n_{\text{ex}}&=n_{su}-KV_{i}^{3}\tau_{Q}^{2},\\ n_{\text{su}}&=\frac{1}{2}-\frac{\alpha}{V_{i}},\end{split}$ (15)
where $K$ and $\alpha$ are fitting parameters. When the quench rate declines and reaches the tuning point, $\tau_{Q}=\tau_{Q}^{S}$, the pre-saturate regime (PSR) appears because the initial potential strength is out of the impulse regime [24]. The defect’s density within the PSR is a linear combination of $\tau_{Q}^{1/2}$ and $\tau_{Q}^{3/2}$:
@@
Figure 4: The defects as a function of quench time $\tau_{Q}$ with system size $L=987$. The final defects are obtained from the linear quench by adjusting the quench rate from slow to very fast. After the fast quench regimes, the fit gives $n_{\text{ex}}\sim\tau_{Q}^{-0.507\pm 0.007}$ (KZ-I) and $n_{\text{ex}}\sim\tau_{Q}^{-0.420\pm 0.008}$ (KZ-II), corresponding to the fast and slow quench rate. The alternation of the KZ exponents reflects the two plateaus of the dynamical exponent $z\approx 1$ and $z\approx 1.388$. At saturate regime (SR), the density of defects is saturated to a plateau at first and then decreases with a scaling $n_{\text{ex}}\sim\tau_{Q}^{2}$. In the pre-saturate regime (PSR), $n_{\text{ex}}$ declines as $1/2-0.52\tau_{Q}^{0.49}$ until $\tau_{Q}>1$.
@@
IV.1.1 Final defects after the one-way quench
@@
The final defects density $n_{\text{ex}}$ as a function of the quench time $\tau_{Q}$ is shown in Fig. 4 from slow to very fast quench rate with the initial potential strength $V_{i}=4$. According to the behavior of the defect density, we indicate the saturated, pre-saturated, and Kibble-Zurek regimes. In the saturated regime (SR), the saturate plateau $n_{\text{su}}=0.3865$ appears when the quench rate is very fast, as the dashed red line. A fitting of $(n_{\text{su}}-n_{\text{ex}})\sim\tau_{Q}^{s}$ yields $s=2.004\pm 0.005$, which is in good agreement with Eq. (15). With the increasing of quench time, the initial potential strength $V_{i}$ is out of the non-adiabatic regime, $V_{i}>V+\hat{\epsilon}$, and quench dynamics turns into pre-saturated regime (PSR). In this case, $n_{\text{ex}}$ diminishes as the expression $1/2-0.52\tau_{Q}^{0.49}$, which is revealed by a fitting of $(1/2-n_{\text{ex}})=\tau_{Q}^{s}$ yields $s=0.490\pm 0.009$. The $\tau_{Q}^{3/2}$ term in Eq. (16) does not appear because of the small fitting parameter $B$. After the quench time gets to $\tau_{Q}\approx 1$, the adiabatic-impulse-adiabatic hypothesis kicks in, and the KZ scaling law can reasonably predict the behavior of defects’ density.
@@
In the KZ regime, the power-law fit and Eq. (13) imply two dynamical exponents $z_{\text{I}}^{D}\approx 0.972\pm 0.027$ and $z_{\text{II}}^{D}\approx 1.381\pm 0.045$ for exact $\nu=1$, corresponding to the moderately fast and slow quench rates. The two values of dynamical exponents match the ones obtained by the finite-size scaling as shown in Fig. (2), $z_{\text{I}}=1$ and $z_{\text{II}}=1.388$, with systematic errors $3\%$ and $0.6\%$ respectively. In the moderately fast quench rate, $z_{\text{I}}$ closely resembles that of the transverse field Ising model. In the slow quench regime, the KZ exponent shifts to $z_{\text{II}}$, which is in good agreement with the known result [65]. ̵Specifically, for the moderately fast quench rate, the impulse regime is far from the critical point but does not extend beyond $V_{i}$ with the dynamical exponent staying at a constant $z(\hat{\epsilon}\gtrsim 1)\approx 1$. In this case, the defect’s density behaves as a power-law scaling $n_{\text{ex}}\sim\tau_{Q}^{-0.5}$. With the quench rate slowing down, the non-adiabatic regime becomes narrow and near the critical point. The dynamical exponent will saturate to a plateau $z(\hat{\epsilon}\rightarrow 0)\approx 1.388$, corresponding to the excitation probability scaling as $n_{\text{ex}}\sim\tau_{Q}^{-0.42}$, which in good agreement with the result in the past research [67]. This variable KZ exponent is similar to that in the random Ising chain affected by the Griffith phase [72, 70]. However, the KZ exponent will be steady because the dynamical exponent saturates to a constant at the critical point instead of diverging.
@@
The influence of the dynamical exponent is also reflected in the non-universal function (14), which discloses a universal behavior of the excitation probability in the impulse region. The defect’s density as a function of scaled time in the non-adiabatic region is shown in Fig. 5(a) and (c) by varying quench time $\tau_{Q}$ belonging to KZ-I and KZ-II, respectively. Different scaling factors, $\tau_{Q}^{0.5}$ and $\tau_{Q}^{0.42}$ corresponding to the inverse of $\hat{n}_{\text{ex}}$, are multiplied in $n_{\text{ex}}$ to disclosed the non-universal function (14) in Fig. 5(c) and (d).
IV.1.2 Final defects with fast quench rate